As many 8cm diameter discs as possible are cut from a sheet of rectangular cardboard measuring 170cm by 90cm. Find the area of the sheet left.
Answer by checkley77(12844)

170/8=21.25
90/8=11.25
21*11=231 disks can be cut from this piece of cardboard.
170*90=15,300 cm^2 total area of the cardboard.
Area of the disks:
231(3.14*4^2)
231*3.14*16
231*50.24=11,605.44 cm^2 total area of all disks.
15,300-11,605.44=3,694.56 cm^2 amount of the sheet left.